OpenStudy (anonymous):

Find the distance CB for the points c(8, -6) and B(6, 7). A) distance = √198 B) distance = √197 C) distance = √172 D) distance = √173

OpenStudy (mertsj):

So now that Parth posted the distance formula for you, why don't you use it?

OpenStudy (mertsj):

Subtract the x coordinates of the points. Subtract the y coordinates of the points. What do you get?

OpenStudy (anonymous):

I've never done this stuff so idk.....

OpenStudy (mertsj):

In the ordered pair, (8,6) which is the x?

OpenStudy (anonymous):

8?

OpenStudy (mertsj):

In the ordered pair (6,7) which is the x?

OpenStudy (anonymous):

6

OpenStudy (mertsj):

Subtract those two numbers. What do you get?

OpenStudy (anonymous):

2

OpenStudy (mertsj):

Remember that number. We'll come back to it. In the ordered pair (8,-6) what is the y?

OpenStudy (anonymous):

-6

OpenStudy (mertsj):

In the ordered pair (6,7) what is the y?

OpenStudy (anonymous):

7

OpenStudy (mertsj):

Subtract those two numbers. What do you get?

OpenStudy (anonymous):

Am I doing -6 - 7 or 7 - 6?

OpenStudy (mertsj):

Doesn't matter if you do -6-7 or 7-(-6)

OpenStudy (mertsj):

Because you are going to square it anyway.

OpenStudy (anonymous):

-6 - 7 = -13 7 - -6 = 1

OpenStudy (mertsj):

7-(-6)=7+6=13

OpenStudy (mertsj):

What was the other number you rememberes?

OpenStudy (mertsj):

remembered?

OpenStudy (anonymous):

8 - 6 = 2

OpenStudy (mertsj):

Square the 2 and square the 13. What do you get?

OpenStudy (mertsj):

\[2^2= ???\] \[13^2=???\]

OpenStudy (anonymous):

2² = 4 & 13²= 169

OpenStudy (mertsj):

Add them together. What do you get?

OpenStudy (anonymous):

do I add those together? Which make 173?

OpenStudy (mertsj):

Take the square root and that is your distance.

OpenStudy (anonymous):

So it's 173?

OpenStudy (mertsj):

That's why I drew you the picture in the question you closed. It's just a special case of the Pythagorean Theorem.

OpenStudy (mertsj):

Did you take the square root?

OpenStudy (anonymous):

Umm, how do you take the square root?

OpenStudy (mertsj):